OpenStudy (anonymous):

how to write two equivalent fraction 10 11

OpenStudy (p0sitr0n):

multiply it by the same number integer at top and bottom per axample 1/2 * 2/2 = 2/4 = 1/2 because 2/2 = 1 and x* 1 = x